OpenStudy (anonymous):

A walrus accelerates from 7.0km/h to 34.5km/h over a distance of 95m. What is the magnitude of the walrus's acceleration?

OpenStudy (anonymous):

$$ Average\ Acceleration = \frac{ Change\ In\ Velocity}{Change\ In\ Time} $$ Simply written as: $$ A_{avg} = \frac{\Delta V}{\Delta t} $$ Do you know what Velocity is equal to?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

As an equation I mean. Do you know the equation for Average velocity?

OpenStudy (anonymous):

Well don't I have to convert 7.0km/h and 34.5km/h to m/s?

OpenStudy (anonymous):

Okay, average velocity is just like acceleration, except with it is your change in *distance* over change in time. $$ V_{avg} = \frac{\Delta x}{\Delta t} $$

OpenStudy (anonymous):

Yes you do

OpenStudy (anonymous):

Okay but I do not have a time in this problem

OpenStudy (anonymous):

Im really new to this stuff and I absolutely can not grasp the stuff. Could you possibly show me exactly what to do